# On the connectedness of the singular set of holomorphic foliations

### v2

@@ Abstract.
Let $\mathscr{F}$ be a codimension 1 singular holomorphic foliation on $\mathbb{P}^{3}$ and let $Z$ be the union of the 1-dimensional connected components of its singular set; Dominique Cerveau asked in 2013 whether $Z$ is necessarily connected. We prove that if a $k$-dimensional holomorphic foliation on a projective manifold whose determinant bundle of the normal sheaf is ample has a singular set $Z$ of dimension $k-1$, then $Z$ is connected. This can be regarded as providing an obstruction for the integrability of singular holomorphic distributions in the spirit of Bott’s work.
@@ Introduction
where $TX$ is the sheaf of holomorphic vector fields on $X$, the tangent sheaf $\mathscr{F}$ is an integrable coherent subsheaf of $TX$ of rank $k$, and the normal sheaf $N_{\mathscr{F}}$ is coherent and torsion free, of rank $n-k;$ this is the codimension of $\mathscr{F}$. The singular scheme, denoted by ${\rm Sing}(\mathscr{F})$, is the locus where the normal sheaf is not locally free. This implies that ${\rm Sing}(\mathscr{F})$ has codimension $\geq 2$. Let $\det(N_{\mathscr{F}})$ denote the determinant bundle of the normal sheaf. We prove:
@@
Let $X$ be a projective manifold of dimension $n$, and $\mathscr{F}$ a singular holomorphic foliation of dimension $k>1$ such that the bundle $\det(N_{\mathscr{F}})$ is ample. Let the singular set be ${\rm Sing}(\mathscr{F})=Z\cup Z_{-}$, where $Z$ has dimension $k-1$ and $Z_{-}$ consists of lower-dimensional components. Then $Z$ is a connected set.
@@
This answers positively the question raised by D. Cerveau in [7] regarding dimension 2 foliations on $\mathbb{P}^{3}$, since the determinant bundle of the normal sheaf of every holomorphic foliation on projective spaces is ample. Cerveau’s problem was motivated by the known models of codimension one foliations classified in degrees 0, 1, and 2.
We recall that in the 1960s (see, for instance, [3]) Raoul Bott investigated global obstructions for the integrability of a subbundle of the tangent bundle of a manifold. Theorem 1 can be regarded as belonging to that frame of thought, as it provides a global obstruction for a singular holomorphic distribution to be integrable. In fact, we know from [5, Theorem 9.5 item 2b.] and [13, Table 1] that there are families of holomorphic distributions of dimension two on $\mathbb{P}^{3}$, whose singular set contains two connected components of dimension 1.
Our proof uses in an essential way the Baum–Bott residues [2]. The idea is as follows. By Baum–Bott, the Chern number $c_{1}(N_{\mathscr{F}})^{n-k+1}[X]$ has a residue localized at each connected component $Z_{1},\dots,Z_{s}$ of the singular set of dimension $k-1$; these residues are elements in $H_{2k-2}(Z_{i})$. Then, the ampleness condition yields, after some work, that the residue at each $Z_{i}$ equals the total Chern number $c_{1}^{n-k+1}(N_{\mathscr{F}})\cap[X]$. Hence $s=1$.
@@
This naturally leads to the following question: What are the topological obstructions for a rank $2$ reflexive sheaf to be the tangent sheaf of a foliation in a projective 3-fold?
Next, we present an application of our theorem that emphasizes its interpretation as a topological obstruction to integrability, by imposing a topological condition on the tangent sheaf and in terms of the number of isolated singularities of the foliation. Hence, this provides an answer to the question above, providing a topological obstruction to integrability.
@@ Acknowledgments
CA and JS would like to thank the University of Bari for its hospitality and support. JS is partially supported by UNAM-PAPIIT grant IN101424. MC is partially supported by the Università degli Studi di Bari and by the PRIN 2022MWPMAB- “Interactions between Geometric Structures and Function Theories” and he is a member of INdAM-GNSAGA; he would like to thank Nigel Hitchin for insightful discussions during his visit to the University of Oxford. MJ is supported by the CNPQ grant number 305601/2022-9, the FAPESP-ANR project number 2021/04065-6.
@@ 1. Proof of Theorem 1
[…]
where all the cohomology and homology groups are with coefficients in $\mathbb{C}$; $P=\,\cap[X]$ is Poincaré duality; the $U_{i}$ are pairwise disjoint smooth regular neighborhoods (as in [15]) of the connected components $Z_{1},\dots,Z_{s}$ of $Z$, $A$ is Alexander duality, all the other homomorphisms are induced by the inclusion. The morphism $\phi$ comes from the following long exact sequence of relative cohomology
@@
First of all, we observe that the previous formula implies that the ampleness of $c_{1}(N_{\mathscr{F}})$ ensures that the component $Z=\bigcup_{i=1}^{s}Z_{i}$ of dimension $k-1$ is non-empty, since $c_{1}(N_{\mathscr{F}})^{n-k+1}\neq 0$, and component $Z_{-}$ of dimension smaller than $k-1$ does not contribute to ${\widehat{\rm Res}}$.
@@
[…]
@@
for $\text{for all }j\neq i\quad\overline{U_{i}}\cap\overline{U_{j}}=\emptyset,\quad(\overline{U_{i}}\setminus Z_{i})\cap Sing(\mathscr{F})=\emptyset.$
@@
Since $Z_{i}$ is connected, we have that $H_{2(k-1)}(Z_{i})\simeq H_{2(k-1)}(U_{i})$ is generated by the fundamental cycle $[Z_{i}]$, it follows that the residue of $\mathscr{F}|_{U_{i}}$ at $Z_{i}$ satisfies:
${\rm Res}(\mathscr{F}|_{U_{i}},Z_{i})=\lambda_{i}[Z_{i}]={\rm Res}(\mathscr{F},Z_{i})\,,$
for some $\lambda_{i}\in\mathbb{C}$. Thus,
$P_{i}^{-1}\big{(}(i_{Z_{i}})_{*}{\rm Res}(\mathscr{F},Z_{i})\big{)}=\lambda_{i}\widehat{[Z_{i}]}\in H^{2(n-k+1)}(X)\;,$
where
$\widehat{[Z_{i}]}:=P^{-1}\big{(}(i_{Z_{i}})_{*}[Z_{i}]\big{)}\,$
the Poincaré dual of $[Z_{i}]$.
Applying $j^{*}_{U_{i}}$ to the previous equation we get the following equality in $\in H_{c}^{2(n-k+1)}(U_{i})$:
(1.3) $j^{*}_{U_{i}}\big{(}\widehat{\rm Res}(\mathscr{F},Z_{i})\big{)}\,=\,\lambda_{i}\cdot j^{*}_{U_{i}}\big{(}\widehat{[Z_{i}]}\big{)}\,=\,j^{*}_{U_{i}}\big{(}c_{1}({\rm det}(N_{\mathscr{F}})\big{)}^{n-k+1}\;,$
@@
where $h:=c_{1}({\mathcal{O}}_{X}(1))$. Hence:
@@
Let ${[Z_{i}]}$ be the fundamental cycle. Substituting in (1.3) we get:
$\lambda_{i}\cdot j^{*}_{U_{i}}\big{(}\widehat{[Z_{i}]}\big{)}=j^{*}_{U_{i}}\big{(}c_{1}(N_{\mathscr{F}})^{n-k+1}\big{)}=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\,,$
Since $k>1$, we can apply $j^{*}_{U_{i}}(h^{k-1})$ in this equation and we get:
$\lambda_{i}\cdot(j^{*}_{U_{i}}\widehat{[Z_{i}]})\smile j^{*}_{U_{i}}(h^{k-1})=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\smile j^{*}_{U_{i}}(h^{k-1})\,,$
That implies:
$\lambda_{i}\cdot j^{*}_{U_{i}}(\widehat{[Z_{i}]}\smile h^{k-1})=r^{n-k+1}j^{*}_{U_{i}}(h^{n})\,.$
Hence
$\widehat{[Z_{i}]}\smile h^{k-1}=P^{-1}[Z_{i}]\smile P^{-1}[W_{i}]=P^{-1}([Z_{i}\cap W_{i}])=d_{i}\cdot P^{-1}[p_{i}],$
where $\{p_{i}\}=Z_{i}\cap W_{i}$ is the intersection point of $Z_{i}$ with a generic complete intersection $W_{i}=H_{1}^{i}\cap\cdots\cap H_{n-k+1}^{i},$ with
$H_{1}^{i},\dots,H_{n-k+1}^{i}\in H^{0}(X,\mathcal{O}_{X}(1)),$
where $d_{i}$ is the multiplicity of this intersection. Notice that $p_{i}\in U_{i}$ can be defined in general by the intersection of hyperplane sections:
$P^{-1}[p_{i}]=[\widetilde{H}_{1}\cap...\cap\widetilde{H}_{n}]=h^{n},$
for generic $\widetilde{H}_{1}\cap...\cap\widetilde{H}_{n}\in H^{0}(X,\mathcal{O}_{X}(1))$. Thus we have
$\lambda_{i}\cdot j^{*}_{U_{i}}(\widehat{[Z_{i}]}\smile h^{k-1})=\lambda_{i}\cdot j^{*}_{U_{i}}(d_{i}\big{(}P^{-1}[p_{i}]\big{)})=r^{n-k+1}j^{*}_{U_{i}}(h^{n}),$
and $j^{*}_{U_{i}}h^{n}=j^{*}_{U_{i}}\big{(}P^{-1}[p_{i}]\big{)}\neq 0$, since $p_{i}\in Z_{i}\subset U_{i}$. Hence,
(1.4) $\lambda_{i}d_{i}=r^{n-k+1}.$
Recall that Baum–Bott says:
$c_{1}(N_{\mathscr{F}})^{n-k+1}=\sum_{i=1}^{s}\lambda_{i}\widehat{[Z_{i}]}.$
Since $c_{1}(N_{\mathscr{F}})^{n-k+1}=r^{n-k+1}h^{n-k+1}$, then by (1.4)
$r^{n-k+1}h^{n-k+1}=\sum_{i=1}^{s}\left(\frac{r^{n-k+1}}{d_{i}}\right)\widehat{[Z_{i}]}\,.$
Applying $h^{k-1}$ we get
$r^{n-k+1}h^{n}$ $=\sum_{i=1}^{s}\left(\frac{r^{n-k+1}}{d_{i}}\right)\cdot\widehat{[Z_{i}]}\smile h^{k-1}$ $=\sum_{i=1}^{s}r^{n-k+1}\frac{d_{i}h^{n}}{d_{i}}$
$=\sum_{i=1}^{s}r^{n-k+1}h^{n},$
since $\widehat{[Z_{i}]}\smile h^{k-1}=d_{i}h^{n}$. Thus
$h^{n}=\sum_{i=1}^{s}h^{n}=sh^{n}$
implies that $(s-1)h^{n}=0$, so $s=1$, since $h$ is ample. Therefore, $Z$ is connected.
@@ 4. Examples and remarks
Let $X$ be the classical Hopf manifold of dimension $n$. This is obtained from $\mathbb{C}^{n}\setminus\{0\}$ dividing it by the group of automorphisms generated by a contraction, say $z\mapsto z/2$. It is clear that $X$ fibers over $\mathbb{P}^{n-1}$ with fiber an elliptic curve, the quotient of $\mathbb{C}^{*}$ by the contraction. So one has an elliptic fibration $X\buildrel{\pi}\over{\to}\mathbb{P}^{n-1}$.
@@
(Foliations on products) Let $\mathscr{F}$ be a holomorphic foliation of dimension $1$ with isolated singularities on a projective manifold $X$ of dimension $n$. If $Y$ is a projective manifold of dimension $k-1$, then we have a foliation $\pi_{2}^{*}\mathscr{F}$ on $X\times Y$ of dimension $k$, with singular set the disjoint union
@@
since $n+k-1>n$. So the determinant of the normal bundle is not ample.
@@
The next example generalizes to high dimensions an example from [4] for $n=3$; the foliation ~$\widetilde{\mathscr{F}}_{3,d}$ described below defines a rigid irreducible component of the space of foliations of codimension one on $\mathbb{P}^{3}$ of degree $d+1$, and ~$\widetilde{\mathscr{F}}_{3,1}$ generates the exceptional component of foliations of degree 2 [4, Cor. 3, p 1007].
@@
Since $[v_{n,d},u_{n,d}]=v_{n,d},$ they are the infinitesimal generators of an action of the affine group, $\mathfrak{aff}(\mathbb{C})\times\mathbb{C}^{n}\longrightarrow\mathbb{C}^{n}.$
@@
[…]
@@
hence, ${\rm Sing}(\widetilde{\mathscr{F}}_{n,d})$ is connected. Notice that the restriction ~$\widetilde{\mathscr{F}}_{n,d}|_{\{Z_{n}=0\}}$ is ~$\widetilde{\mathscr{F}}_{n-1,d}$, and
@@
Let $\mathscr{F}$ be a holomorphic foliation on $\mathbb{P}^{n}$ of dimension $k>1$, with $Z\subset{\rm Sing}(\mathscr{F})$ the non-empty component of dimension $k-1$. We can write
@@
$[Z_{i}\cap Z_{k}]=(d_{j}h^{n-k+1})\smile(d_{k}h^{n-k+1})=d_{j}d_{k}h^{2(n-k+1)}\in H^{2(n-k+1)}(\mathbb{P}^{n}),$
with $d_{j}=\deg(Z_{j})$ and $d_{k}=\deg(Z_{k})$. Thus, $[Z_{i}\cap Z_{k}]\neq 0$ whenever $n\leq 2k-2$, implying that $Z$ is connected in this case. Consequently, the statement of Theorem 1 is not immediate in the case where $n>2k-2$. Recently, in [10], it was shown that $\mathrm{dim}({\rm Sing}(\mathscr{F}))\geq k-1$, whenever $n\geq 2k$, meaning that $Z_{-}$ is empty.
@@
shows that the section $\nu_{\mathscr{F}}$ corresponds to a twisted $(n-k)$-form
@@
with values in $\det(N_{\mathscr{F}})$. A holomorphic foliation $\mathscr{F}$ of dimension $k$ on $\mathbb{P}^{n}$ can be described by a twisted $(n-k)$-form
@@
where $Z_{ij}$ are its irreducible components, for $i=1,\dots,s$. Theorem 1 implies the component of dimension $k-1$ of the singular set of the foliation $Z$ is connected. That is, $Z=Z_{i}$, for all $i$, and by Baum–Bott

### v3

@@ Abstract.
Let $\mathscr{F}$ be a singular holomorphic foliation of dimension $k>1$ in a projective $n$-manifold $X$. Assume that the normal sheaf of $\mathscr{F}$ is ample (as is always the case when $X=\mathbb{P}^{n}$) and the singular set $\mbox{Sing}(\mathscr{F})$ has dimension $\leq k-1$. We show that the union of the irreducible components in $\mbox{Sing}(\mathscr{F})$ of dimension exactly k1k-1 is necessarily connected. Consequently, this yields a Bott-type topological obstruction to the integrability of singular holomorphic distributions, echoing Bott’s vanishing theorem, and it answers a question by D. Cerveau for codimension one foliations on $\mathbb{P}^{3}$.
@@ Introduction
In the 1960s, Raoul Bott [3] demonstrated that topological constraints can obstruct the existence of globally integrable foliated structures. In the same vein, we establish a global obstruction to integrability for singular holomorphic distributions. Indeed, there exist holomorphic distributions of dimension $k>1$ on $\mathbb{P}^{n}$, for all $n\geq 3$, whose singular set contains several connected components of dimension k1k-1 (see Remark 4.5). Our main Theorem precludes integrability for all such distributions.
@@
where $TX$ is the sheaf of holomorphic vector fields on $X$, the tangent sheaf $\mathscr{F}$ is an integrable coherent subsheaf of $TX$ of rank $k$, and the normal sheaf $N_{\mathscr{F}}$ is coherent and torsion free, of rank nk;n-k; this is the codimension of $\mathscr{F}$. The singular scheme, denoted by ${\rm Sing}(\mathscr{F})$, is the locus where the normal sheaf is not locally free. This implies that ${\rm Sing}(\mathscr{F})$ has codimension $\geq 2$. Let $\det(N_{\mathscr{F}})$ denote the determinant bundle of the normal sheaf. We prove:
@@
Let $X$ be a projective manifold of dimension $n$, and $\mathscr{F}$ a singular holomorphic foliation of dimension $k>1$ such that the bundle $\det(N_{\mathscr{F}})$ is ample. Let the singular set be ${\rm Sing}(\mathscr{F})=Z\cup Z_{-}$, where $Z$ has dimension k1k-1 and $Z_{-}$ consists of lower-dimensional components. Then $Z$ is a connected set.
@@
Theorem 1 answers positively the question posed by D. Cerveau in [7] regarding dimension 2 foliations on $\mathbb{P}^{3}$, since the determinant bundle of the normal sheaf of every holomorphic foliation on projective spaces is ample. Cerveau’s problem was motivated by the known models of codimension one foliations classified in degrees 0, 1, and 2.
This falls within the Bott phenomenon of global obstructions to integrability: whereas Bott’s theorem compels the vanishing of invariant polynomials in the Chern classes of the normal bundle in degrees exceeding the codimension for integrable subbundles, our connectedness Theorem 1 furnishes a topological obstruction to integrability for singular distributions under an ampleness hypothesis. We know from [5, Theorem 9.5 item 2b.] and [13, Table 1] that there are families of holomorphic distributions of dimension two on $\mathbb{P}^{3}$, whose singular set contains two connected components of dimension 1.
Our proof uses in an essential way the Baum–Bott residues [2]. The idea is as follows. By Baum–Bott, the Chern number $c_{1}(N_{\mathscr{F}})^{n-k+1}[X]$ has a residue localized at each connected component $Z_{1},\dots,Z_{s}$ of the singular set of dimension k1k-1; these residues are elements in $H_{2k-2}(Z_{i})$. Then, the ampleness condition yields, after some work, that the residue at each $Z_{i}$ equals the total Chern number $c_{1}^{n-k+1}(N_{\mathscr{F}})\cap[X]$. Hence $s=1$.
@@
We show in Section 4 that the ampleness hypothesis on $\det(N_{\mathscr{F}})$ is indispensable for Theorem 1. Moreover, when $X=\mathbb{P}^{n}$ and $n\leq 2(k-1)$, any two irreducible projective subvarieties $V_{1},V_{2}\subset\mathbb{P}^{n}$ of dimension k1k-1 must meet. Indeed, in $A^{*}(\mathbb{P}^{n})=\mathbb{Z}[h]/(h^{n+1})$ we have
$V_{1}\cap V_{2}]=\deg(V_{1})\deg(V_{2})h^{\,2n-2k+2}\neq 0.$
Consequently, the (k1)(k-1)-dimensional components of ${\rm Sing}(\mathscr{F})$ cannot be pairwise disjoint, and the conclusion of Theorem 1 is automatic in this range. However, this phenomenon does not hold in general. For example, consider $X=\mathbb{P}^{1}\times\mathbb{P}^{5}$ and subvarieties $W_{1}=\{p\}\times\mathbb{P}^{4}$ and $W_{2}=\{q\}\times\mathbb{P}^{4}$, where $p\neq q$. Here, $n=6$ and $k=4$, so n=2(k1)n=2(k-1), yet $W_{1}$ and $W_{2}$ are clearly disjoint.
@@
This naturally leads to the following question:
What are the topological obstructions for a rank $2$ reflexive sheaf to be the tangent sheaf of a foliation in a projective 3-fold?
Answering the preceding question in the setting of projective threefolds, we show that a natural topological constraint on the tangent sheaf—together with the ampleness of the normal determinant—yields a global obstruction to integrability, detected by the number of isolated singularities of the foliation.
@@ Acknowledgments
CA and JS gratefully acknowledge the University of Bari for its hospitality and support through UNIBA Visiting Professor Program. JS is partially supported by UNAM-PAPIIT grant IN101424. MC is partially supported by the Università degli Studi di Bari and by the PRIN 2022MWPMAB- “Interactions between Geometric Structures and Function Theories” and he is a member of INdAM-GNSAGA; he would like to thank Nigel Hitchin for insightful discussions during his visit to the University of Oxford. MJ is supported by the CNPQ grant number 305601/2022-9, the FAPESP-ANR project number 2021/04065-6.
@@ 1. Proof of Theorem 1
$H^{2(n-k+1)}(X,X\setminus Z)$$H^{2(n-k+1)}(X)$$H^{2(n-k+1)}(U_{i},U_{i}\setminus Z_{i})$$H_{c}^{2(n-k+1)}(U_{i})$$H_{2k-2}(Z)$$H_{2k-2}(X)$$H_{2k-2}(Z_{i})$$H_{2k-2}(U_{i})$$\scriptstyle{\phi}$$\scriptstyle{A}$$\scriptstyle{j^{*}_{U_{i}}}$$\scriptstyle{P}$$\scriptstyle{j^{*}_{U_{i}}}$$\scriptstyle{\phi_{i}}$$\scriptstyle{A_{i}}$$\scriptstyle{P_{i}}$$\scriptstyle{(i_{Z})_{*}}$$\scriptstyle{P^{-1}}$$\scriptstyle{(i_{Z_{i}})_{*}}$$\scriptstyle{(j_{Z_{i}})_{*}}$$\scriptstyle{(i_{Z_{i}})_{*}}$$\scriptstyle{A_{i}^{-1}}$$\scriptstyle{(j_{U_{i}})_{*}}$$\scriptstyle{P_{i}^{-1}}$
where all the cohomology and homology groups are with coefficients in $\mathbb{C}$; $P=\,\cap[X]$ is Poincaré duality; the $U_{i}$ are pairwise disjoint smooth regular neighborhoods (as in [15]) of the connected components $Z_{1},\dots,Z_{s}$ of $Z$, AA is Alexander duality, all the other homomorphisms are induced by the inclusion. The morphism $\phi$ comes from the following long exact sequence of relative cohomology
@@
First of all, we observe that the previous formula implies that the ampleness of $c_{1}(N_{\mathscr{F}})$ ensures that the component $Z=\bigcup_{i=1}^{s}Z_{i}$ of dimension k1k-1 is non-empty, since $c_{1}(N_{\mathscr{F}})^{n-k+1}\neq 0$, and component $Z_{-}$ of dimension smaller than k1k-1 does not contribute to $^{\widehat{\rm Res}}$.
@@
${Z_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{Z_{i}}}$${U_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho_{i}}$${Z_{i}}$
@@
$\text{for all }j\neq i\quad\overline{U_{i}}\cap\overline{U_{j}}=\emptyset,\quad(\overline{U_{i}}\setminus Z_{i})\cap Sing(\mathscr{F})=\emptyset.$
@@
Write $Z_{i}$ as the union of its irreducible components of dimension k1k-1:
$Z_{i}=\bigcup_{\alpha=1}^{m_{i}}W_{i\alpha}\qquad(\mathrm{dim}W_{i\alpha}=k-1).$
The localized residue in $U_{i}$ then reads
$\widehat{\mathrm{Res}(\mathscr{F}|_{U_{i}},Z_{i})}=\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,[\widehat{W_{i\alpha}}]\ \in H_{c}^{2(n-k+1)}(U_{i}).$
Applying $j^{*}_{U_{i}}$ we get the following equality in $H_{c}^{2(n-k+1)}(U_{i})$:
(1.3) $j^{*}_{U_{i}}\big(\widehat{\rm Res}(\mathscr{F},Z_{i})\big)\,=\,\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\big)\,=\,j^{*}_{U_{i}}\big(c_{1}({\rm det}(N_{\mathscr{F}})\big)^{n-k+1}\big),$
@@
where $h:=c_{1}({\mathcal{O}}_{X}(1))$. Hence
@@
Since $k>1$, we can apply $j^{*}_{U_{i}}(h^{k-1})$ in (1.3) and we get
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\!\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n-k+1})\smile j^{*}_{U_{i}}(h^{k-1})\,$,
which implies
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=r^{n-k+1}j^{*}_{U_{i}}(h^{n})\,.$
For each $\alpha$ we have
$\widehat{W_{i\alpha}}]\smile h^{k-1}=P^{-1}[W_{i\alpha}]\smile P^{-1}[W_{i}]=P^{-1}\big([W_{i\alpha}\cap W_{i}]\big)=d_{i\alpha}\,P^{-1}[p_{i\alpha}],$
where $\{p_{i\alpha}\}=W_{i\alpha}\cap W_{i}$ is the (scheme-theoretic) intersection point of $W_{i\alpha}$ with a generic complete intersection
$W_{i}=H_{1}^{i}\cap\cdots\cap H_{n-k+1}^{i},\qquad H_{1}^{i},\dots,H_{n-k+1}^{i}\in H^{0}(X,\mathcal{O}_{X}(1)),$
and
$d_{i\alpha}=\operatorname{length}(W_{i\alpha}\cap W_{i})=(h^{k-1}\!\cdot W_{i\alpha})>0.$
Moreover, $p_{i\alpha}\in U_{i}$ can be defined by hyperplane sections through $p_{i\alpha}$:
$P^{-1}[p_{i\alpha}]=[\,\widetilde{H}^{(i,\alpha)}_{1}\cap\cdots\cap\widetilde{H}^{(i,\alpha)}_{n}\,]=h^{n},\qquad\widetilde{H}^{(i,\alpha)}_{\ell}\in H^{0}(X,\mathcal{O}_{X}(1))\ \text{generic}.$
Thus
$\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big([\widehat{W_{i\alpha}}]\smile h^{k-1}\big)=\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,j^{*}_{U_{i}}\big(d_{i\alpha}P^{-1}[p_{i\alpha}]\big)=r^{n-k+1}\,j^{*}_{U_{i}}(h^{n}),$
and $j^{*}_{U_{i}}(h^{n})=j^{*}_{U_{i}}\big(P^{-1}[p_{i\alpha}]\big)\neq 0$ because $p_{i\alpha}\in Z_{i}\subset U_{i}$. Hence,
(1.4) $\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,d_{i\alpha}=r^{\,n-k+1}.$
Baum–Bott in cohomology gives
$c_{1}(N_{\mathscr{F}})^{\,n-k+1}=\sum_{i=1}^{s}\widehat{\mathrm{Res}}(\mathscr{F},Z_{i})=\sum_{i=1}^{s}\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}\,[\widehat{W_{i\alpha}}$].
Cupping with $h^{k-1}$ and using $\widehat{W_{i\alpha}}]\smile h^{k-1}=d_{i\alpha}\,h^{n}$ and (1.4), we obtain
$r^{\,n-k+1}h^{n}=c_{1}(N_{\mathscr{F}})^{\,n-k+1}\smile h^{k-1}=\sum_{i=1}^{s}\left(\sum_{\alpha=1}^{m_{i}}\lambda_{i\alpha}d_{i\alpha}\,h^{n}\right)=\sum_{i=1}^{s}r^{\,n-k+1}h^{n}.$
Hence $h^{n}=s\,h^{n}$, and since $h$ is ample, $h^{n}\neq 0$; therefore $s=1$ and $Z$ is connected.
@@ 4. Examples and remarks
Let $X$ be the classical Hopf manifold of dimension $n$. This is obtained from $\mathbb{C}^{n}\setminus\{0\}$ dividing it by the group of automorphisms generated by a contraction, say $z\mapsto z/2$. It is clear that $X$ fibers over $\mathbb{P}^{n-1}$ with fiber an elliptic curve, the quotient of $\mathbb{C}^{*}$ by the contraction. So one has an elliptic fibration $X\mathrel{\mathop{\kern 0.0pt\to}\limits^{{\pi}}}\mathbb{P}^{n-1}$.
@@
(Foliations on products) Let $\mathscr{F}$ be a holomorphic foliation of dimension $1$ with isolated singularities on a projective manifold $X$ of dimension $n$. If $Y$ is a projective manifold of dimension k1k-1, then we have a foliation $\pi_{2}^{*}\mathscr{F}$ on $X\times Y$ of dimension $k$, with singular set the disjoint union
@@
since n+k1>nn+k-1>n. So the determinant of the normal bundle is not ample.
@@
The next example generalizes to high dimensions an example from [4] for $n=3$; the foliation $\widetilde{\mathscr{F}}_{3,d}$ described below defines a rigid irreducible component of the space of foliations of codimension one on $\mathbb{P}^{3}$ of degree $d+1$, and $\widetilde{\mathscr{F}}_{3,1}$ generates the exceptional component of foliations of degree 2 [4, Cor. 3, p 1007].
@@
Since $v_{n,d},u_{n,d}]=v_{n,d},$ they are the infinitesimal generators of an action of the affine group, $\mathfrak{aff}(\mathbb{C})\times\mathbb{C}^{n}\longrightarrow\mathbb{C}^{n}.$
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widetilde{\mathscr{F}}_{n,d}={\mathcal{O}}_{\mathbb{P}^{n}}\oplus{\mathcal{O}}_{\mathbb{P}^{n}}(1-d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{u}_{n,d}\oplus\widetilde{v}_{n,d}}$${T\mathbb{P}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\widetilde{\mathscr{F}}_{n,d}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
@@
hence, ${\rm Sing}(\widetilde{\mathscr{F}}_{n,d})$ is connected. Notice that the restriction $\widetilde{\mathscr{F}}_{n,d}|_{\{Z_{n}=0\}}$ is $\widetilde{\mathscr{F}}_{n-1,d}$, and
@@
Let $\mathscr{F}$ be a holomorphic foliation on $\mathbb{P}^{n}$ of dimension $k>1$, with $Z\subset{\rm Sing}(\mathscr{F})$ the non-empty component of dimension k1k-1. We can write
@@
$Z_{i}\cap Z_{k}]=(d_{j}h^{n-k+1})\smile(d_{k}h^{n-k+1})=d_{j}d_{k}h^{2(n-k+1)}\in H^{2(n-k+1)}(\mathbb{P}^{n}),$
with $d_{j}=\deg(Z_{j})$ and $d_{k}=\deg(Z_{k})$. Thus, $Z_{i}\cap Z_{k}]\neq 0$ whenever $n\leq 2k-2$, implying that $Z$ is connected in this case. Consequently, the statement of Theorem 1 is not immediate in the case where n>2k2n>2k-2. Recently, in [10], it was shown that $\mathrm{dim}({\rm Sing}(\mathscr{F}))\geq k-1$, whenever $n\geq 2k$, meaning that $Z_{-}$ is empty.
@@
shows that the section $\nu_{\mathscr{F}}$ corresponds to a twisted (nk)(n-k)-form
@@
with values in $\det(N_{\mathscr{F}})$. A holomorphic foliation $\mathscr{F}$ of dimension $k$ on $\mathbb{P}^{n}$ can be described by a twisted (nk)(n-k)-form
@@
where $Z_{ij}$ are its irreducible components, for $i=1,\dots,s$. Theorem 1 implies the component of dimension k1k-1 of the singular set of the foliation $Z$ is connected. That is, $Z=Z_{i}$, for all $i$, and by Baum–Bott